OpenStudy (anonymous):

y varies directly as the square of x. When x = 3, y = 54. Find y when x = 8 please help

jimthompson5910 (jim_thompson5910):

"y varies directly as the square of x" translates to y = kx^2

jimthompson5910 (jim_thompson5910):

When x = 3, y = 54, so plug these values in to get y = kx^2 54 = k*3^2 Then solve for k 54 = k*9 54/9 = k 6 = k k = 6 So we get the equation y = 6x^2

jimthompson5910 (jim_thompson5910):

Your last step is to plug in x = 8 into y = 6x^2 and evaluate to find y.

OpenStudy (anonymous):

thank you :) .. so what are these kind of problems callled

jimthompson5910 (jim_thompson5910):

variation problems, in this case, it's a direct variation problem

jimthompson5910 (jim_thompson5910):

you'll also come across inverse variation problems, but that's for another problem

OpenStudy (anonymous):

thank you so much #i appreciate u helping me

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

thank you # algebra 2 is killing lolz

jimthompson5910 (jim_thompson5910):

just keep at it and you'll get it eventually

OpenStudy (anonymous):

thx:)